D D $24 MC ATC S 21 19 S 160 240 320 8 10 At price of $24, each firm makes a profit of $ 16 24 30 72

D D MC ATC $24 S 21 19 S 160 240 320 8 10 In the start, market equilibrium is units 80 160 240 320

D D MC ATC $24 S 21 19 S 160 240 320 8 10 In the start, the number of firms in this industry is 20 24 30 40

D S MC ATC D 16 $12 8 240 360 720 6 12 At price of $8, each firm makes a _ profit. 48 -24 0 24 48

D $24 MC ATC S 21 19 S 160 240 320 8 10 At price of $24, ATC is $: 10 19 21 24

D MC ATC DO 16 $12 8 240 360 720 6 12 D falls to D, which brings market quantity to: 0 240 360 720

D S MC ATC 16 $12 8 240 360 720 6 12 In the end, each firm manufactures_units. 6 8 12 16

D s s MC ATC DI 16 $12 8 240 360 720 6 12 At price of $8, each unit is made for a_profit. 8

D S $ MC ATC D 16 $12 8 240 360 720 6 12 In the end, this market is served by _ firms 20 30 40 60

D S S MC ATC 16 D $12 8 240 360 720 6 12 At price of $8, ATC is $: 4 12 16

D S MC ATC DO 16 $12 00 240 360 720 6 12 The first market equilibrium is units OO 240 360 720

D S S MC ATC D 16 $12 8 240 360 720 6 12 in the end, the final market price is $: 6 08 12 16

D D MC ATC $24 S 21 19 s 160 240 320 8 10 In the end, the number of firms in this industry is: 20 24 30 40

D D $24 MC ATC S 21 19 S 160 240 320 CO 10 D surges to D, which causes market equilibrium quantity to now be. 80 160 240 3

D S MC ATC D 16 $12 co 240 360 720 6 12 in the end, the final market quantity is: 120 240 360 720
Answer #1

1) Answer is C. $30.

Profit = quantity (price - ATC) = 10 (24-21 ) = 10× 3 = 30.

2) answer is B. 160 units.

Market equilibrium is where market supply and market demand intersects.

3) Answer is A. 20

Total units = 160.

Units by one firm = 8

Number of firms = 160/8 = 20.

4) answer is A. -48

Profit = quantity (price - ATC) = 6(8-16) = 6× -8 = -48.

5) answer is C. $21

#Please rate positively...thank you